OpenStudy (anonymous):

48 is 12% of what number

OpenStudy (anonymous):

X*12/100=48 X=400

OpenStudy (anonymous):

0.12*X=48 48/0.12 = x; x=400

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

aer you sure because there is no x

OpenStudy (anonymous):

i took the unknown no. as X

OpenStudy (anonymous):

ah ok

OpenStudy (anonymous):

you can check 400*12/100=48